OpenStudy (anonymous):

What is the solution set of 4x^2 + 8x - 5 = 0?

OpenStudy (anonymous):

usa test prep?

OpenStudy (anonymous):

Algebra Final Review.

OpenStudy (anonymous):

oh yeah i saw that problem while working yesterday, let me see if i remember

OpenStudy (anonymous):

:) there ya go

OpenStudy (anonymous):

use it as a guide

OpenStudy (anonymous):

OK, let's see what I get...

OpenStudy (anonymous):

Wait, I don't really see anything I can factor out... Hmm...

OpenStudy (anonymous):

oh i see, you can use the quadratic formula

OpenStudy (anonymous):

Oh, OK. I'll try that.

OpenStudy (anonymous):

that should give you 1/2 i believe

OpenStudy (anonymous):

You're right!

OpenStudy (anonymous):

:) glad i can help

OpenStudy (anonymous):

But wait, there might be a second answer.

OpenStudy (anonymous):

Hence "solution set".

OpenStudy (anonymous):

did you get 1/2?

OpenStudy (anonymous):

Yes, one of the answers is 1/2. I tested it out and it works.

OpenStudy (anonymous):

oh yeah and the 2nd answer is 5/2

OpenStudy (anonymous):

if you're wondering how i got it heres the work! 4x^2+8x-5=0 => 4x^2 - 2x + 10x -5 = 0 => 2x(2x - 1 ) + 5(2x-1) = 0 => (2x-1)(2x+5) = 0 => x = 1/2 ; x = -5/2 Ans

OpenStudy (anonymous):

So the answers are 1/2 and -5/2? OK, got it. Thanks!

OpenStudy (anonymous):

no problem